OpenStudy (anonymous):

1/7 = 3/ 4x-1 solve for x Will Fan!

OpenStudy (anonymous):

HELP please

OpenStudy (anonymous):

I am lost please

OpenStudy (anonymous):

add one the the left side

OpenStudy (anonymous):

to the*

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

they are subtracting

OpenStudy (anonymous):

\[1/7=3/4x-1\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

you have to get x on one side by itself so you have to add one to both sides

OpenStudy (anonymous):

I am confused

OpenStudy (anonymous):

We are solving for X so you have to get rid of the -1 by adding 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5x

OpenStudy (anonymous):

oh no 4x

OpenStudy (anonymous):

\[8/7=3/4x\]

OpenStudy (anonymous):

thanks for your help but i am going to ask my teacher

OpenStudy (anonymous):

lol okay